How to solve {y^2 - (3/24)^2} {y^2-(1/24)^2} = 5/864?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Write,
{y^2 - (3/24)^2} {y^2-(1/24)^2} as =>
{y^2 - (1/24)^2 – 8/(24)^2} {y^2-(1/24)^2} = 5/864
let {y^2 - (1/24)^2} = xthen,(x -8/(24)^2) x = 5/864Now solve this quadratic in x and then equate it to general y form.